# Prompting Explicit and Implicit Knowledge for Multi-hop Question Answering Based on Human Reading Process thanks: This paper has been accepted at LREC-COLING 2024

### 01 Mar 2024 v2

@@ 4.3.   Main Results and Analysis
Table 1: Results on the blind test set of HotpotQA in the distractor setting. “-” denotes the case where no results are available. ${}^{{\dagger}}$ denotes that we implement the codes. “Ans” represents the metrics for answer; “Sup” denotes the metrics for supporting facts; “Joint” is the joint metrics that combine the evaluation of answer spans and supporting facts. HotpotQA Leaderboard: https://hotpotqa.github.io/.
We compare our PEI model with other published baselines on the test set of HotpotQA in the distractor setting, including the baseline model of HotpotQA (Yang et al., 2018), SOTA model on the leaderboard Zhang et al. (2023), iCAP Wang et al. (2022) and PCL Deng et al. (2022) which we inspired by, and other baselines. As shown in Table 1, our proposed PEI model demonstrates superior performance across all evaluation metrics compared to all other baselines (except Beam Retrieval), and achieves comparable performance with Beam Retrieval Zhang et al. (2023) on the HotpotQA dataset, highlighting the significant progress made by PEI for multi-hop QA.
Specifically, our PEI achieves an improvement of $0.20$ in answer EM, $0.28$ in answer F1 score, and $0.30$ in join F1 score when compared to Beam Retrieval. Conversely, Beam Retrieval shows an improvement of $1.22$ in supporting EM, $0.28$ in supporting F1 score, and $0.62$ in join EM when compared to PEI.
@@ 4.4.   Evaluation of Robustness
Evaluation on Other Multi-hop Datasets. To verify the generalization, we evaluate PEI model on 2WikiMultihopQA and MuSiQue. In Table 2, the results show that our PEI model surpasses all comparison baselines in terms of both EM and F1 metrics. Notably, despite both PEI and iCAP employ the same encoder-decoder BART model, PEI achieves a substantial improvement of $4.52/26.66$ in the answer EM/F1 score compared to iCAP on the 2WikiMultihopQA dataset. Furthermore, our model exhibits superior performance compared to PCL, yielding improvements of $1.29/1.14$ and $0.69/0.51$ in the answer EM/F1 score on 2WikiMultihopQA and MuSiQue, respectively.
@@
Table 2: Results of our proposed PEI compared to PCL, HGN and iCAP on 2WikiMultihopQA and MuSiQue multi-hop QA test set.
Evaluation on Sub-question Dataset. To evaluate the PEI model’s efficacy in the multi-hop reasoning process, specifically in composing answers from solved sub-questions, we conducted an evaluation on sub-question QA dataset (Tang et al., 2021). This dataset consists of 1000 multi-hop questions $q$ along with their corresponding sub-questions $q_{sub1}$ and $q_{sub2}$. Table 3 shows that our PEI model achieves a $97.62\%$ success rate in correctly answering the parent multi-hop question $q$ when both sub-questions $q_{sub1}$ and $q_{sub2}$ are answered correctly  [^2] The calculation process: $49.2/(49.2+1.2)=97.62\%$. This highlights the proficiency of PEI in retaining acquired knowledge through the integration of explicit and implicit knowledge, surpassing other multi-hop QA models. However, it is noteworthy that PEI also demonstrates a significant likelihood (36.55%) of correctly answering the parent multi-hop question even when only one of the sub-questions is answered accurately  [^3] The calculation process: $(7.1+22.1)/(49.2+7.1+22.1+1.5)=36.55\%$. To further illustrate the sub-question dependent success rate of various multi-hop QA models, Figure 3 shows that these models exhibit a high probability (exceeding $20\%$) of correctly answering parent multi-hop questions even when only one sub-question is answered correctly. This indicates that the utilization of potentially unsound reasoning shortcuts in predicting answers is a prevalent and challenging phenomenon in multi-hop QA tasks.
@@
Table 3: Categorical EM statistics (%) of sub-question evaluation for five multi-hop QA models. $c/w$ denotes that the question is answered correctly/wrongly. $sub1$ denotes the first sub-question and $sub2$ is the second sub-question of corresponding question $q$. For example, the first four rows show the percentage of multi-hop questions that can be correctly answered.
Figure 3: The success rate (%) of five multi-hop QA models. $sub1$ denotes the first sub-question and $sub2$ is the second sub-question of corresponding question $q$.
@@ 4.5.   Ablation Studies
Effect of Foundation PLMs. To assess the effects of foundation PLMs, we compare PEI with PCL and HGN based on the same data and backbone. As shown in Table 5, PEI consistently outperforms PCL and HGN across all metrics. This indicates the effectiveness and robustness of PEI across PLMs. Moreover, PEI with ALBERT achieves an improvement of $0.62/0.23$ in Ans/Sup F1 compared to PEI with RoBERTa. ALBERT outperforms RoBERTa on GLUE benchmark using a single-model setup  [^4] https://github.com/google-research/albert. These results confirm that adopting a more competent foundation PLM can improve the performance of our model.
@@
Table 5: Results with different PLMs on the development set of HotpotQA. RoBERTa, ELECTRA and ALBERT denote that we use RoBERTa-large, ELECTRA-large and ALBERT-xxlargev2 as the PLM respectively

### 04 Mar 2024 v3

@@ 4.3.   Main Results and Analysis
Table 1: Results on the blind test set of HotpotQA in the distractor setting. “-” denotes the case where no results are available. ${}^{{\dagger}}$ denotes that we implement the codes. Other results are derived from (Deng et al., 2022) and (Zhang et al., 2023). “Ans” represents the metrics for answer; “Sup” denotes the metrics for supporting facts; “Joint” is the joint metrics that combine the evaluation of answer spans and supporting facts.
Initially, we evaluate our proposed PEI model on the test set of HotpotQA in the distractor setting comparing with peer-reviewed baselines, including the baseline model of HotpotQA (Yang et al., 2018), SOTA model on the leaderboard Zhang et al. (2023), iCAP Wang et al. (2022) and PCL Deng et al. (2022) which we inspired by, and other baselines. Illustrated in Table 1, our proposed PEI model demonstrates superior performance across all evaluation metrics compared to all other baselines (except Beam Retrieval), and achieves comparable performance with Beam Retrieval Zhang et al. (2023) on the HotpotQA dataset, highlighting the significant progress made by PEI for multi-hop QA.
More specifically, our PEI achieves an improvement of $0.20$ in answer EM, $0.28$ in answer F1 score, and $0.30$ in join F1 score when compared to Beam Retrieval. Conversely, Beam Retrieval shows an improvement of $1.22$ in supporting EM, $0.28$ in supporting F1 score, and $0.62$ in join EM when compared to PEI.
@@ 4.4.   Evaluation of Robustness
Evaluation on Other Multi-hop Datasets. To assess the generalization, we evaluate PEI model on 2WikiMultihopQA and MuSiQue. In Table 2, the results show that our PEI model surpasses all comparison baselines in terms of both EM and F1 metrics. Notably, despite both PEI and iCAP employ the same encoder-decoder BART model, PEI achieves a substantial improvement of $4.52/26.66$ in the answer EM/F1 score compared to iCAP on the 2WikiMultihopQA dataset. Furthermore, our model exhibits superior performance compared to PCL, yielding improvements of $1.29/1.14$ and $0.69/0.51$ in the answer EM/F1 score on 2WikiMultihopQA and MuSiQue, respectively.
@@
Table 2: Results of our proposed PEI compared to PCL, HGN (Deng et al., 2022) and iCAP (Wang et al., 2022) on 2WikiMultihopQA and MuSiQue multi-hop QA test set. “-” denotes the case where no results are available.
Evaluation on Sub-question Dataset. To evaluate the PEI model’s efficacy in the multi-hop reasoning process, specifically in composing answers from solved sub-questions, we conducted an evaluation on sub-question QA dataset (Tang et al., 2021). The parent questions are denoted as $q$ along with their corresponding sub-questions $q_{sub1}$ and $q_{sub2}$. Table 3 shows that our PEI model achieves a $97.62\%$ success rate in correctly answering the parent multi-hop question $q$ when both sub-questions $q_{sub1}$ and $q_{sub2}$ are answered correctly  [^2] The calculation process: $49.2/(49.2+1.2)=97.62\%$. This highlights the proficiency of PEI in retaining acquired knowledge through the integration of explicit and implicit knowledge, surpassing other multi-hop QA models. However, it is noteworthy that PEI also demonstrates a significant likelihood (36.55%) of correctly answering the parent multi-hop question even when only one of the sub-questions is answered accurately  [^3] The calculation process: $(7.1+22.1)/(49.2+7.1+22.1+1.5)=36.55\%$. To further illustrate the sub-question dependent success rate of various multi-hop QA models, Figure 3 shows that these models exhibit a high likelihood of correctly answering parent multi-hop questions even when only one sub-question is answered correctly. Therefore, in multi-hop QA, it is a prevalent and unsolved problem that models unexpectedly leverage unreliable reasoning shortcuts for answer predictions, which we do not expect that to happen Deng et al. (2022).
@@
Table 3: Results on sub-question dataset. $c/w$ denotes that the question is answered correctly/wrongly. $sub1$ denotes the first sub-question and $sub2$ is the second sub-question of corresponding question $q$. The results of DFGN, DecompRC, HGN and PCL are derived from (Deng et al., 2022).
Figure 3: The success rate (%) of five multi-hop QA models. $sub1$ denotes the first sub-question and $sub2$ is the second sub-question of corresponding question $q$. The results of DFGN, DecompRC, HGN and PCL are from (Deng et al., 2022).
@@ 4.5.   Ablation Studies
Effect of Foundation PLMs. To assess the effects of foundation PLMs, we compare PEI with PCL and HGN based on same experimental context including data and foundation models. Illustrated in Table 5, PEI consistently outperforms PCL and HGN across all metrics. This indicates the effectiveness and robustness of PEI across PLMs. Moreover, PEI with ALBERT achieves an improvement of $0.62/0.23$ in Ans/Sup F1 compared to PEI with RoBERTa. ALBERT outperforms RoBERTa on GLUE benchmark using a single-model setup  [^4] https://github.com/google-research/albert. These results confirm that adopting a more competent foundation PLM can improve the performance of our model.
@@
Table 5: Results with different PLMs on the development set of HotpotQA. RoBERTa, ELECTRA and ALBERT represent RoBERTa-large (#params: 355M), ELECTRA-large (#params: 335M) and ALBERT-xxlargev2 (#params: 223M) as the PLM respectively. The results of HGN and PCL are derived from (Deng et al., 2022).